OpenStudy (anonymous):

[3.05] MA.7.A.3.4 What operation must be done to both sides of the equation to solve p minus negative three and seven-ninths equals five-sixths? Add five-sixths Subtract three and seven-ninths Subtract five-sixths Add three and seven-ninths

OpenStudy (anonymous):

Okay. Hmm. The equation looks like: \[p - (-3) + \frac{ 7 }{ 9 } = \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

The multiple choice in my opinion doesn't really make sense, but the answer I would choose is, "Add three and seven-ninths"

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @miamoore